# Chinese vs. World Bank Development Projects: Insights from Earth Observation and Computer Vision on Wealth Gains in Africa, 2002-2013

### v1

@@ 5.3 Insights about Aid Assignment Mechanisms
To complement our image–adjusted IPW estimates, we also estimate a two–way fixed–effects model that relies only on within–neighborhood changes in exposure among treatment switchers:
@@
where $Y_{it}$ is the IWI measured one period after commitment, $A_{it}$ indicates sector–funder treatment in period $t$, $\alpha_{i}$ are neighborhood (DHS–cluster) fixed effects, and $\lambda_{t}$ are period fixed effects. Standard errors are clustered at the ADM2 level. This design differences out all time–invariant local characteristics (baseline wealth, geography, built environment together) and thus reduces reliance on selection‑on‑observables (similar to Bitzer and Gören (2018), but using fixed effects units 451 times more granular in terms of area (552 km2 vs. 6.72 km2); identification comes from units that switch treatment status over time, so estimates are more conservative, less precise, use only variation from units where treatment status switches—something that would not be possible without satellite-based neighborhood-level imputations from Pettersson et al. (2023).
Results (see Appendix) are directionally consistent with the main analysis but markedly attenuated: across most sectors the estimated ATT remains positive for both funders, yet magnitudes are substantially smaller and confidence intervals wider, reflecting the limited number of switchers. The broad China >> World Bank pattern largely persists, though the gap narrows. We interpret these within–unit estimates as a conservative robustness check that supports the main conclusion—aid generally raises local wealth—while highlighting that effects are smaller when identified purely from computer-vision-powered within‑neighborhood variation.
@@ 6 Discussion
Researchers continually seek to improve the causal understanding of the impact of local development projects on individuals’ health and living conditions (Imbens and Rubin, 2015). In the case of World Bank and Chinese, there is a lack of localized tabular data on the geographical and historical maps that officials used to strategically decide on where to allocate their projects across Africa in the past two decades (Bedi et al., 2007a). Our study relies on a novel EO-ML method that proxies these maps with the help of satellite image Jerzak (2023b, a). Another innovation of our study is that we use a novel subnational data on health and living conditions recently produced by computer scientist Pettersson et al. (2023). Combining these two innovations, this study analyzes the effect of locally targeted World Bank and Chinese development projects. The study encompasses a sector-specific analysis, yielding findings that are roughly comparable to those of the subnational studies. Our use of nearly 10,000 neighborhoods as potential treated and control observations for each funder and sector analysis enables individual-sector-level analysis, making visible significant sector heterogeneity. These data and methodological innovations advance the aid effectiveness debate by addressing some of the challenges facing subnational development studies. Next, we discuss potential implication of our results and methodology.
Hypothesizing Potential Mechanisms Driving the Results: Why Beneficial Effects?
A key finding of our analysis is that, on average, Chinese aid programs demonstrate a larger positive impact on our machine-learning-derived wealth index than World Bank programs across nearly all comparable sectors (see Figure 7). This divergence in effect size can likely be attributed to the differing mechanisms through which each donor’s project portfolio and implementation style influence local economies. The literature widely notes that Chinese aid is heavily concentrated in physical infrastructure sectors, such as transportation, energy, and construction (Brautigam, 2009; Dreher et al., 2018).
@@
In contrast, the World Bank’s projects yield smaller, though still broadly positive and statistically significant, effects on the IWI. The Bank’s official strategy involves working through recipient government systems to build long-term capacity, a process codified by the Paris Declaration on Aid Effectiveness (OECD, 2005). The causal mechanisms for projects in sectors like “Government and Civil Society” or “Other Social Infrastructure” are therefore more indirect. They aim to improve the quality and accountability of public services, effects that are often slow to materialize and may not translate into measurable household wealth within our study’s timeframe. Even for a sector with a direct link to the IWI, like “Water Supply and Sanitation,” the World Bank’s process-oriented approach may be slower than China’s more direct implementation (Isaksson and Kotsadam, 2018). The weaker performance of World Bank aid in “Industry, Mining, Construction” and “Agriculture,” as found in our analysis, further suggests its model may be less suited to generating the kind of immediate, capital-intensive impact that our wealth measure captures effectively (Ssozi et al., 2019).
@@
Data and Methodological Contributions.
Pettersson et al. (2023) wealth estimates overcome measurement limitations in nightlight data used to proxy subnational wealth, overcome the lack of panel data in Demographic and Health Surveys, and enable the use of fixed effects at the ADM2 (county, district, city) level to control for unobserved, time-invariant confounders in subnational regions.
@@
Data resolution is a concern in studies using satellite methods and data. If the resolution of satellite images is not sufficiently high, the algorithm may miss confounding factors. The multi-resolution nature of our data raises several challenges to statistical inference (see Meng (2014)), including issues related to mixing data measured at multiple resolutions and to estimating an estimand at a finer-grained resolution (5 km2 neighborhoods) than many of the covariates upon which it is based.
To ensure the use of nationally representative samples in each country included in each funder/sector analysis, we use locations of neighborhoods surveyed in one round of Demographic and Health Surveys as units of analysis. We use Pettersson et al. (2023)’s wealth estimates for those locations over time. This is a non-random sample, since all these countries have a relationship with the U.S. Future research could use wealth estimates to create broader samples of neighborhoods across the African continent, including in countries never surveyed by DHS, or could include all human settlement areas across Africa instead of a sample. With the possible dismantling of funds to USAID, and its detrimental effect on future DHS data collection efforts, EO-ML methods such as Pettersson et al. (2023) will be even more critical for aid evaluation.

### v2

@@ 5.4 Robustness and Alternative Identification Strategies
To reduce reliance on selection on observables, we estimate a two-way fixed-effects model that uses only within-neighborhood changes in exposure among treatment switchers:
@@
where $Y_{it}$ is IWI (measured one period after commitment), $A_{it}$ indicates sector–funder treatment in period $t$, $\alpha_{i}$ are neighborhood fixed effects, and $\lambda_{t}$ are period fixed effects. Standard errors are clustered at the ADM2 level. Across Chinese projects, point estimates are directionally consistent with the IPW results (>0>0) but attenuated in magnitude and less precise (see Figure 10), reflecting the limited number of switchers. World Bank effects more often switch signs in this unit-level fixed effects analysis. For the World Bank sector with the most consistently negative point estimates (Government and Civil Society (150)), the corresponding fixed-effects estimate is closer to zero and imprecise (-0.05, SE = 0.11). We interpret these results cautiously, pending a more rigorous investigation into switcher dynamics and within-sector heterogeneity that would complicate causal interpretation.
Figure 10: Unit-level fixed effects results.
Robustness to lagged nightlights in the adjustment set.
Because lagged per-capita nightlights are persistent and also enter the IWI construction pipeline, we re-estimate the main ViT M+X+FE specifications after removing lagged nightlights from $\mathbf{X}_{i(t-1)}$. Estimates remain qualitatively consistent: across the 2525{} funder–sector cells, the median absolute ATE change is 2.05 IWI points and the maximum absolute change is 7.38. The correlation between baseline and no-nightlights ATEs is 0.85.
Robustness to neighborhood size.
To assess sensitivity to the 6.7 km $\times$ 6.7 km outcome definition—and the possibility that DHS displacement moves some clusters toward neighborhood edges—we construct a coarser IWI outcome by averaging the four adjacent 6.7 km tiles around each cluster (effective 13.4 km neighborhoods) and re-estimate the main models. ATEs show high but slightly lower correlation with the baseline specification (correlation: 0.60) compared to the nightlight robustness; the qualitative ordering of effects across funders and sectors is broadly unchanged.
Robustness to treatment definition.
Finally, we tighten the treatment definition by excluding ADM2-wide assignments that arise from imprecisely geocoded (precision-3) projects and restricting treatment to precision-1 and precision-2 locations. Figure A.I.7 shows that effects remain broadly consistent with the main analysis; baseline and strict-precision ATEs are correlated (correlation: 0.61); see the Appendix. That said, World Development estimates are notably higher in magnitude in this regime.
Overall, these robustness tests suggest that conclusions from the main analysis about the relative ranking of sector effects and the relative effects of Chinese vs. World Bank programs are not generally sensitive to modeling choices about nightlights, outcome resolution, or treatment-definition rules. Choice of treatment and especially identification strategy have greatest consequence for the magnitude and precision of sector-specific estimates. Broadly speaking, across identification strategy and treatment/outcome definition, Chinese programs are still estimated here to have greater effects on living conditions.
@@ 6 Discussion
Researchers continually seek to improve the causal understanding of the impact of local development projects on individuals’ health and living conditions (Imbens and Rubin, 2015). In the case of the World Bank and China, there is a lack of localized tabular data on the geographical and historical maps that officials used to strategically decide on where to allocate their projects across Africa in the past two decades (Bedi et al., 2007a). Our study applies and extends a recently proposed EO–ML method that proxies these maps with the help of satellite images Jerzak (2023b, a). The contribution here is less about inventing a new estimator and more about demonstrating, in a large‑scale applied setting, how image‑based confounder adjustment changes estimated impacts and what it reveals about assignment mechanisms. Another innovation here is that we use a novel subnational dataset on health and living conditions recently produced by Pettersson et al. (2023). Combining these two advances, this study analyzes the effect of locally targeted World Bank and Chinese development projects. The study includes a sector-specific analysis, yielding findings roughly comparable to those of the subnational studies. Our use of nearly 10,000 neighborhoods as potential treated and control observations for each funder and sector analysis enables individual-sector-level analysis, making visible significant sector heterogeneity. Our findings are also robust to dropping lagged nightlights from the covariate set and to substituting alternative proxies for baseline economic conditions, alleviating concerns about mechanically re‑using nightlights in both outcome construction and confounder adjustment. These results advance the aid effectiveness debate by addressing some of the challenges facing subnational development studies. Next, we discuss potential implications.
Hypothesizing Potential Mechanisms Driving the Results: Why and Where Beneficial Effects?
A finding of this analysis is that, on average, Chinese aid programs demonstrate a larger positive impact on our machine-learning-derived wealth index than World Bank programs across nearly many comparable sectors (see Figure 7). This divergence in effect size can likely be attributed to the differing mechanisms through which each donor’s project portfolio and implementation style influence local economies. The literature notes that Chinese aid is heavily concentrated in physical infrastructure sectors, such as transportation, energy, and construction (Brautigam, 2009; Dreher et al., 2018).
@@
In contrast, the World Bank’s projects yield smaller, though still broadly positive and often statistically significant, effects on the IWI. The Bank’s official strategy involves working through recipient government systems to build long-term capacity, a process codified by the Paris Declaration on Aid Effectiveness (OECD, 2005). The causal mechanisms for projects in sectors like “Government and Civil Society” or “Other Social Infrastructure” are therefore more indirect. They aim to improve the quality and accountability of public services, effects that are often slow to materialize and may not translate into measurable household wealth within our study’s timeframe. Even for a sector with a direct link to the IWI, like “Water Supply and Sanitation,” the World Bank’s process-oriented approach may be slower than China’s more direct implementation (Isaksson and Kotsadam, 2018). The weaker performance of World Bank aid in “Industry, Mining, Construction” and “Agriculture,” as found in our analysis, further suggests its model may be less suited to generating the kind of immediate, capital-intensive impact that our wealth measure captures effectively (Ssozi et al., 2019).
@@
Additionally, Pettersson et al. (2023) wealth estimates overcome measurement limitations in nightlight data used to proxy subnational wealth, overcome the lack of panel data in Demographic and Health Surveys, and enable the use of fixed effects at the ADM2 (county, district, city) level to control for unobserved, time-invariant confounders in subnational regions.
@@
Data resolution is a concern in studies using satellite methods and data. If the resolution of satellite images is not sufficiently high, the algorithm may miss confounding factors. The multi-resolution nature of our data raises several challenges to statistical inference (see Meng (2014)), including issues related to mixing data measured at multiple resolutions and to estimating an estimand at a finer-grained resolution (6.7 km neighborhoods) than many of the covariates upon which it is based. To ensure the use of nationally representative samples in each country included in each funder/sector analysis, we use locations of neighborhoods surveyed in one round of Demographic and Health Surveys as units of analysis. We use Pettersson et al. (2023)’s wealth estimates for those locations over time. This is a non-random sample, since all these countries have a relationship with the U.S. Future research could use wealth estimates to create broader samples of neighborhoods across the African continent, including in countries never surveyed by DHS, or could include all human settlement areas across Africa instead of a sample. With the possible dismantling of funds to USAID, and its detrimental effect on future DHS data collection efforts, EO-ML methods such as Pettersson et al. (2023) will be even more critical for aid evaluation.
@@
Data and code availability.
All code to reproduce the analyses in this paper, along with replication datasets derived from publicly available sources, will be made available at
and